You know that you will need $30,000 for a down payment on a house in 10 years. If your account earns 6% compounded monthly, how much would you need to deposit now to reach your goal?
Round your answer to the nearest cent.

Answer by stanbon(75887)

You know that you will need $30,000 for a down payment on a house in 10 years. If your account earns 6% compounded monthly, how much would you need to deposit now to reach your goal?
-----
A(t) = P*(1+(r/n))^(nt)
------
30,000 = P*(1+0.06/12)^(12*10)
------
30,000 = P(1.005)^120
----
30,000 = P(1.8194)
---
P = $16,488.98
---------
Cheers,
Stan H.
---------